# The Misclassification Likelihood Matrix: Some Classes Are More Likely To Be Misclassified Than Others

### v1

@@ 1 Introduction
Researchers have studied and addressed the varying hazards of misclassifications in different domains. [KW06] explore the relationship between diversity measures and ensemble accuracy in classifier ensembles. They discuss the importance of considering the costs of different types of errors and how they can be incorporated into the design and evaluation of classifier ensembles. [MS20] propose a framework for consistent cost-sensitive learning with nonlinear loss functions. They address the issue of varying misclassification costs and present algorithms that can effectively handle such scenarios, enabling the learning of classifiers that are sensitive to the relative hazards of different types of errors.
Ensuring the reliability of automated decision-making systems is crucial in high-stakes domains where errors can lead to severe consequences [AOS16]. Machine learning models, despite their impressive performance, are susceptible to making inaccurate predictions when presented with data that deviates from the training distribution, a problem known as distribution shift [HBM21, QCSSL09].
Distribution shift occurs when the training data distribution differs from the deployment data distribution, leading to a decline in model performance [QCSSL09, HD19]. It can manifest as covariate shift, concept drift, or domain shift [MTRAR12, Shi00, GŽB14, PGLC15]. To quantify and mitigate distribution shift, researchers employ statistical divergence measures [KL51, GBR12], domain adaptation techniques [WD18, SKM07, PY09, GUA16], and adaptive algorithms [LLD18, BGdCÁF06].
@@ 2.1 Softmax Output And Environmental Aspects
The softmax output of a classifier represents the predicted probabilities for each class. In an $n$-class classification problem, the softmax function takes the logits (raw outputs) of the classifier and transforms them into a probability distribution over the $n$ classes [GBC16]. The softmax output is a vector of length $n$, where each element represents the predicted probability of the input belonging to the corresponding class. The probabilities in the softmax output sum up to 1, indicating the relative confidence of the classifier in each class prediction. The class with the highest probability is typically considered the predicted class. The softmax output provides a more informative and interpretable representation of the classifier’s predictions compared to the raw logits.
@@
When making a final decision, it is common practice to select only the class with the highest probability, effectively discarding the remaining $(n-1)/n$ of the softmax output [Gal16]. This approach, while straightforward and widely used, raises concerns about the inefficient use of the model’s output and its potential environmental implications.
@@
By considering only the class prediction, we are essentially discarding a significant portion of the information generated by the model. In a scenario with a large number of classes, this can lead to a substantial waste of computational resources and energy [SGM19]. The discarded softmax probabilities, which contain valuable information about the model’s uncertainty and the relationships between classes, are effectively consigned to a digital landfill, contributing to the growing problem of electronic waste and energy consumption in the field of machine learning [SDSE20].
@@
The environmental impact of this practice becomes more pronounced when considering the increasing scale and complexity of modern machine learning models. As the number of classes grows and models become more sophisticated, the amount of discarded information also increases, leading to a larger digital carbon footprint [LLSD19]. This is particularly concerning given the rapid growth of machine learning applications in various domains, from image and speech recognition to natural language processing and autonomous systems [TGLM20].
@@
Certain approaches lend themselves well to make better use of the entire softmax output. One such approach is to incorporate uncertainty estimation techniques, such as Monte Carlo dropout [GG16] or ensemble methods [LPB17], which leverage the full probability distribution to quantify the model’s confidence in its predictions. By considering the uncertainty information, systems can make more informed decisions and avoid discarding potentially useful information [KG17].
@@
Another approach is to develop more efficient and environmentally-friendly machine learning techniques, such as model compression [HMD15], quantization [GKD21], and energy-aware training [GMRRG21]. These methods aim to reduce the computational and memory requirements of models while maintaining their performance, thereby reducing the environmental impact of machine learning systems.
@@
Furthermore, there is a growing awareness of the need for sustainable AI practices and the development of green AI frameworks [SDSE20]. These initiatives focus on designing and implementing machine learning systems that prioritize energy efficiency, resource conservation, and environmental sustainability. By adopting sustainable AI practices, we can mitigate the environmental impact of discarded softmax outputs and other inefficiencies in machine learning pipelines.
@@ 2.2 Using All Softmax Output Predictions
Utilizing the entire set of softmax prediction probabilities, instead of solely depending on the maximum output, has been widely investigated to improve the safety, robustness, and trustworthiness of machine learning models. By considering the complete distribution of class predictions, more reliable and informative prediction pipelines can be developed, enabling techniques such as uncertainty quantification, anomaly detection, and robustness against adversarial attacks [KG17, LPB17, HG17, GSS14, SZS13, MMS17, GRCVDM18].
The softmax probabilities provide valuable information for interpretability and explainability of the model’s decisions, facilitating human-machine collaboration and trust [RSG16, DVK17]. The importance of leveraging the entire softmax distribution extends to various safety-critical domains, such as autonomous vehicles, medical diagnosis, and financial risk assessment, where the consequences of incorrect predictions can be severe [MKG18, LAA17]. By considering the full distribution of predictions, more informed and reliable decisions can be made, reducing the risk of catastrophic failures.
@@ 3.1 Neural Networks
We study not only the highest probability score (prediction), but the entire neural network output vector $\mathbf{p}=(p_{1},p_{2},\dots,p_{K})$ where $\sum p_{i}=1$, representing a probability distribution obtained by normalizing the logit vector $\mathbf{z}=(z_{1},z_{2},\dots,z_{K})$ through the softmax function, $p_{i}=\text{softmax}(z_{i})=e^{z_{i}}/\sum_{j=1}^{K}e^{z_{j}}$. For example, given $\mathbf{p}=[0.01,0.01,0.01,0.01,0.9,0.01,0.01,0.01,0.01,0.01]$, the predicted class is ’4’, corresponding to the highest value at index five, reflecting the confidence of the prediction for each class from ’0’ to ’9’. The logits, representing log-likelihoods of class memberships, are related to probabilities by $z_{i}=\log(p_{i}/(1-p_{i}))$ where $z_{i}$ is the logit for class $i$, and $p_{i}$ is the probability of the input belonging to class $i$[GBC16, Bis06].
@@ 3.2 Perturbed MNIST Dataset
To test our trained model, we augment the MNIST testing dataset, consisting of 10,000 images with 10 classes of approximately 1,000 examples each. We create 12 perturbations - Brightness, Contrast, Defocus Blur, Fog, Frost, Gaussian Noise, Impulse Noise, Motion Blur, Pixelation, Shot Noise, Snow and Zoom Blur - following prior work by [HD19], at 10 intensity levels, generating a dataset of 1,2010,000 images. We modify perturbation function parameters by trial and error such that increased perturbation levels on the testing dataset cause network predictive accuracies to decrease linearly.
@@ 3.3 Misclassification Likelihood Matrix
Figure 2: Perturbed MNIST testing dataset MLM, where dark blue indicates greater, and lighter yellow lesser likelihood likelihood of misclassification.

### v2

# Top row: MNIST digit 5 subject to 10 increasing levels of pixelation Bottom row: Average softmax for all MNIST digit 5 classes across 12 perturbation types, subject to 10 increasing perturbation levels

@@ 1 Introduction
Researchers have studied and addressed the varying hazards of misclassifications in different domains. [kuncheva2006measures] explore the relationship between diversity measures and ensemble accuracy in classifier ensembles. They discuss the importance of considering the costs of different types of errors and how they can be incorporated into the design and evaluation of classifier ensembles. [mozannar2020consistent] propose a framework for consistent cost-sensitive learning with nonlinear loss functions. They address the issue of varying misclassification costs and present algorithms that can effectively handle such scenarios, enabling the learning of classifiers that are sensitive to the relative hazards of different types of errors.
Ensuring the reliability of automated decision-making systems is crucial in high-stakes domains where errors can lead to severe consequences [amodei2016concrete]. Machine learning models, despite their impressive performance, are susceptible to making inaccurate predictions when presented with data that deviates from the training distribution, a problem known as distribution shift [hendrycks2021many, quinonero2009dataset].
Distribution shift occurs when the training data distribution differs from the deployment data distribution, leading to a decline in model performance [quinonero2009dataset, hendrycks2019benchmarking]. It can manifest as covariate shift, concept drift, or domain shift [moreno2012unifying, shimodaira2000improving, gama2014survey, patel2015visual]. To quantify and mitigate distribution shift, researchers employ statistical divergence measures [kullback1951information, gretton2012kernel], domain adaptation techniques [wang2018deep, sugiyama2007covariate, pan2009survey, ganin2016domain], and adaptive algorithms [lu2018learning, baena2006early].
@@ 2.1 Softmax Output And Environmental Aspects
The softmax output of a classifier represents the predicted probabilities for each class. In an $n$-class classification problem, the softmax function takes the logits (raw outputs) of the classifier and transforms them into a probability distribution over the $n$ classes [goodfellow2016deep]. The softmax output is a vector of length $n$, where each element represents the predicted probability of the input belonging to the corresponding class. The probabilities in the softmax output sum up to 1, indicating the relative confidence of the classifier in each class prediction. The class with the highest probability is typically considered the predicted class. The softmax output provides a more informative and interpretable representation of the classifier’s predictions compared to the raw logits.
@@
When making a final decision, it is common practice to select only the class with the highest probability, effectively discarding the remaining $(n-1)/n$ of the softmax output [gal2016uncertainty]. This approach, while straightforward and widely used, raises concerns about the inefficient use of the model’s output and its potential environmental implications.
@@
By considering only the class prediction, we are essentially discarding a significant portion of the information generated by the model. In a scenario with a large number of classes, this can lead to a substantial waste of computational resources and energy [strubell2019energy]. The discarded softmax probabilities, which contain valuable information about the model’s uncertainty and the relationships between classes, are effectively consigned to a digital landfill, contributing to the growing problem of electronic waste and energy consumption in the field of machine learning [schwartz2020green].
@@
The environmental impact of this practice becomes more pronounced when considering the increasing scale and complexity of modern machine learning models. As the number of classes grows and models become more sophisticated, the amount of discarded information also increases, leading to a larger digital carbon footprint [lacoste2019quantifying]. This is particularly concerning given the rapid growth of machine learning applications in various domains, from image and speech recognition to natural language processing and autonomous systems [thompson2020computational].
@@
Certain approaches lend themselves well to make better use of the entire softmax output. One such approach is to incorporate uncertainty estimation techniques, such as Monte Carlo dropout [gal2016dropout] or ensemble methods [lakshminarayanan2017simple], which leverage the full probability distribution to quantify the model’s confidence in its predictions. By considering the uncertainty information, systems can make more informed decisions and avoid discarding potentially useful information [kendall2017uncertainties].
@@
Another approach is to develop more efficient and environmentally-friendly machine learning techniques, such as model compression [han2015deep], quantization [gholami2021survey], and energy-aware training [garcia2021estimation]. These methods aim to reduce the computational and memory requirements of models while maintaining their performance, thereby reducing the environmental impact of machine learning systems.
@@
Furthermore, there is a growing awareness of the need for sustainable AI practices and the development of green AI frameworks [schwartz2020green]. These initiatives focus on designing and implementing machine learning systems that prioritize energy efficiency, resource conservation, and environmental sustainability. By adopting sustainable AI practices, we can mitigate the environmental impact of discarded softmax outputs and other inefficiencies in machine learning pipelines.
@@ 2.2 Using All Softmax Output Predictions
Utilizing the entire set of softmax prediction probabilities, instead of solely depending on the maximum output, has been widely investigated to improve the safety, robustness, and trustworthiness of machine learning models. By considering the complete distribution of class predictions, more reliable and informative prediction pipelines can be developed, enabling techniques such as uncertainty quantification, anomaly detection, and robustness against adversarial attacks [kendall2017uncertainties, lakshminarayanan2017simple, hendrycks17baseline, goodfellow2014explaining, szegedy2013intriguing, madry2017towards, guo2018countering].
The softmax probabilities provide valuable information for interpretability and explainability of the model’s decisions, facilitating human-machine collaboration and trust [ribeiro2016should, doshi2017towards]. The importance of leveraging the entire softmax distribution extends to various safety-critical domains, such as autonomous vehicles, medical diagnosis, and financial risk assessment, where the consequences of incorrect predictions can be severe [michelmore2018evaluating, leibig2017leveraging]. By considering the full distribution of predictions, more informed and reliable decisions can be made, reducing the risk of catastrophic failures.
@@ 3.1 Neural Networks
We study not only the highest probability score (prediction), but the entire neural network output vector $\mathbf{p}=(p_{1},p_{2},\dots,p_{K})$ where $\sum p_{i}=1$, representing a probability distribution obtained by normalizing the logit vector $\mathbf{z}=(z_{1},z_{2},\dots,z_{K})$ through the softmax function, $p_{i}=\text{softmax}(z_{i})=e^{z_{i}}/\sum_{j=1}^{K}e^{z_{j}}$. For example, given $\mathbf{p}=[0.01,0.01,0.01,0.01,0.9,0.01,0.01,0.01,0.01,0.01]$, the predicted class is ’4’, corresponding to the highest value at index five, reflecting the confidence of the prediction for each class from ’0’ to ’9’. The logits, representing log-likelihoods of class memberships, are related to probabilities by $z_{i}=\log(p_{i}/(1-p_{i}))$ where $z_{i}$ is the logit for class $i$, and $p_{i}$ is the probability of the input belonging to class $i$[goodfellow2016deep, bishop2006pattern].
@@ 3.2 Perturbed MNIST Dataset
To test our trained model, we augment the MNIST testing dataset, consisting of 10,000 images with 10 classes of approximately 1,000 examples each. We create 12 perturbations - Brightness, Contrast, Defocus Blur, Fog, Frost, Gaussian Noise, Impulse Noise, Motion Blur, Pixelation, Shot Noise, Snow and Zoom Blur - following prior work by [hendrycks2019benchmarking], at 10 intensity levels, generating a dataset of 1,2010,000 images. We modify perturbation function parameters by trial and error such that increased perturbation levels on the testing dataset cause network predictive accuracies to decrease linearly.
@@ 3.3 Misclassification Likelihood Matrix
Figure 2: Perturbed MNIST testing dataset MLM, where dark blue indicates greater, and lighter yellow lower likelihood of misclassification.